OpenStudy (anonymous):

what are the next 5 numbers in this sequence? 23/24,11/12,21/24,5/6,19/24,3/4

OpenStudy (anonymous):

So what you want to do is look at the numbers you have. Looking at the first two. \[(23\div24) - (11\div12) = (1/24)\] Now lets look at the next set. \[(11\div12) - (21\div24) = (1/24)\] So now that you can see the difference between each subsequent fraction is (1/24) you can solve for the next 5!

OpenStudy (anonymous):

still a bit confused. so 3/4 - X= 1/24????

OpenStudy (anonymous):

correct, and just solve for x